Circular No. 8931

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NR TRIANGULI AUSTRALIS = NOVA TRIANGULI AUSTRALIS 2008

Nicholas J. Brown, Quinns Rocks, Western Australia, reports his discovery of a possible nova (mag 9.2) on two T-Max 400 films taken on Apr. 1.73 UT with a 135-mm f/2 camera lens, giving the variable’s position approximately as α = 16^{h}18^{m}47^{s}, δ = −60^{o}27′57′′ (equinox 2000.0; uncertainty estimated as ∼ 10′′); Brown adds that no star is visible at this position on his film from Mar. 16.76 (limiting mag 11.5), and nothing is visible to mag ∼ 14 in this position on the Digitized Sky Survey. A catalogued star of mag 13.5–14.0 is visible to the northwest of the variable. Brown’s visual observation of the variable on Apr. 2.74 yields mag ≈ 9.0; he also estimated visual mag 10.1 on Apr. 4.695. Following posting on the Central Bureau’s unconfirmed-objects webpage, A. Amorim, Florianopolis, Brazil, reports the following visual magnitude estimates for the variable: Apr. 3.158, 9.3; 4.097, 9.9. W. Liller, Viña del Mar, Chile, reports that a Technical Pan photo (limiting mag ∼ 11.0) taken with an 85-mm camera lens on Mar. 31.33 through an orange filter shows a faint but definite image of the apparent nova at magnitude ≈ 10.5; nothing brighter than mag ∼ 11 appears on a photo taken by Liller on Mar. 15. E. Guido and G. Sostero write that their remote imaging with a 0.31-m f/9 reflector of the Remote Astronomical Society (RAS) located near Moorook, Australia, on Apr. 4.86 yields the following precise position for the presumed nova: α = 16^{h}18^{m}48^{s}.21, δ = −60^{o}27′48′′.9 (equinox 2000.0; UCAC-2 catalogue reference stars). C. Jacques, Belo Horizonte, Brazil, also remotely obtained a 10-s unfiltered CCD image on Apr. 4.875 with an RAS 25-cm f/6 reflector at the Moorook Observatory, yielding mag 10.2 and position end figures 48^{s}.20, 49′′.1. Jacques adds that comparison with a European Southern Observatory red image taken on 1984 Apr. 6 shows nothing at this position (estimated limiting magnitude 19.2), the nearest visible star having position end figures 47^{s}.99, 45′′.8. N. N. Samus, Institute of Astronomy, Russian Academy of Sciences, informs us that this nova is being assigned the variable-star designation NR TrA.

COMET C/2007 W1 (BOATTINI) Visual total-magnitude and coma-diameter estimates: Mar. 6.08 UT,

12.7, 2′.2 (J. J. Gonzalez, Leon, Spain, 0.20-m reflector); 29.07, 10.3, 8′ (M. Goiato, Araçatuba, Brazil, 0.22-m refl.); 31.53, 10.2, 3′.2 (K. Yoshimoto, Yamaguchi, Japan, 0.25-m reflector); Apr. 2.09, 9.5, 6′ (Gonzalez, 25×100 binoculars); 3.98, 8.9, 10′ (Gonzalez, Asturias, Spain, 10×50 binoculars).

2008 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green